Just 57
QUESTION 56 You deposit$ 100 in a savings account that pays 12%interest compounded annually. How much would you have in your account at the end of 5 years? O$160 $166.16 O $168.41 $172.35 O $176.23 QUESTION 57 In the above question, if the bank pays 12% interest compounded qua erly. How much would you have in your account at the end of 5 years? $165.12 $173.16 $180.61 $181.35 $182.45
Answer #1

rate = 12% / 4 = 3%

Number of periods = 5 * 4 = 20

Future value = present value * ( 1 + r)n

Future value = 100 * ( 1 + 0.03)20

Future value = 100 * 1.806111

Future value = $180.61